Ticket Number: HEA-NOC/20050928-6 Ticket Status: SUPERCEDE
Ticket Type: unscheduled Resolver:
Ticket Opened: 20050928 13:51 UTC+1 Problem Start: 20050928 13:43 UTC+1
Ticket Closed: 20060208 15:38 UTC Problem End: 20060108 15:30 UTC
Superceded by: 20060131-7
Both circuits to DIT have reported transitions, and are currently down.
Some traffic to and from DIT may be interupted.
20050928 13:50 BB DIT have spotted possible CPU congestion on their access
router and are investigating further.
20051128 13:53 BN These transitions would appear to be reoccuring. DIT
are currently investigating.
20051206 13:18 DC Circuits transitioning again. DIT are investigating.
20060206 13:41 OB Both circuits have begun transitioning frequently again.
Dearbhla & Eoin are investigating
20060207 10:59 OB Circuits have stopped flapping, however substantial
delay variation is evident
20060208 15:37 BN This issue is being tracked with more detail in Ticket
20060131-7. Any further updates will be tracked by that ticket.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20050928-6